How to express the algebraic snetence if you worked 38 hours a week, you earn 8.50 an hour for monday through friday and 12.75 on saturday and you earned 340 that particular week, how many hours did you work on saturday?
Answer by jrfrunner(365)

let W=hours worked during the week MOnday thru Friday
let S=hours worked Saturday
--
given
hours equation 1: W+S=38 and thus W=38-S
money earned equation 2: 8.50*W+12.75*S=340
--
multiply equation 2 by 100 to eliminate decimals
850*W+1275*S=34000
--
substitute W=38-S into the above equation
850(38-S)+1275*S=34000
32300-850*S+1275*S=34000
425*S=1700
S=4, substitute into W=38-S=38-4=34
--
validate
8.50*W+12.75*S=340
8.50*34+12.75*4=289+51=340 checks out
--
answer: 34 hours during the week and 4 hours on Saturday